OpenStudy (anonymous):

Which of the following is the slope between the two points (0, -4) and (-2, 5)? -nine halves two ninths -two ninths nine halves

OpenStudy (nubeer):

m = (y2-y1)/(x2-x1) apply this formula.. plug in the values u have you will get the slope

OpenStudy (anonymous):

\[\frac{ 5-(-4) }{ -2-0 }\] \[\frac{ 9 }{ -2 }\]